1) 3/16= (?)% x 3/4


2) 5.5/20%=

3) slove for N. N/3=6/9
I need help in trying to solve these probelms, thank you.

Answer by checkley75(3666)

3/16=X*3/4 DIVIDE BOTH SIDES BY 3/4
X=(3/16)/(3/4) NOW INVERT THE DENOMINATOR & MULTIPLY
X=3/16*4/3
X=12/48
X=1/4 OR .25 OR 25% ANSWER.
----------------------------------------
5.5/20%
5.5/.2
2.75 ANSWER.
--------------------------------
N/3=6/9 CROSS MULTIPLY
9N=18
N=18/9
N=2 ANSWER.